5. Use the frequency distribution to the​ right, which shows the number of voters​ (in millions)...

5.

Use the frequency distribution to the​ right, which shows the number of voters​ (in millions) according to​ age, to find the probability that a voter chosen at random is in the given age range.

not between 18 to 20 years old.

The probability is __.

​(Round to three decimal places as​ needed.)

Ages of voters

FrequencyFrequency

18 to 20

5.9

21 to 24

11.9

25 to 34

21.7

35 to 44

24.9

45 to 64

51.8

65 and over

25.7

6. stem and leaf

1| 3 4 5 7 8 8 ​ Key:​1|5=15

2| 0 1 2 2 3 4 5 7 8 8 9

3| 0 0 1 1 1 2 2 3 3 4 4 5 5 7 7 7 8 9

4| 0 0 1 2 2 3 4 5 5 5 5 6 7 8 8 9

5| 0 2 3 5 5 7 8

6| 1 3 6 7

7|

8| 7

​(a) If a team is selected at​ random, find the probability the team scored at least

33 touchdowns ___.

​(Round to three decimal places as​ needed.)

​(b) If a team is selected at​ random, find the probability the team scored between 40 and 49 touchdowns inclusive ___.

​(Round to three decimal places as​ needed.)

​(c) If a team is selected at​ random, find the probability the team scored more than 75 touchdowns ___.

​(Round to three decimal places as​ needed.)

Are any of these events​ unusual? Select all the unusual events below.

A.Scoring between 40 and 49 touchdowns inclusive is unusual.

B.Scoring at least 33 touchdowns is unusual.

C.Scoring more than 75 touchdowns is unusual.

D. None of the events are unusual.
Answer #1

\\ 5) \ 1-\frac{5.9}{5.9+11.9+21.7+24.9+51.8+25.7}=1-0.04158\approx 0.958\\ \\ \\ \\ 6)\ a)\ \frac{22}{63}=0.349\\\\ b) \ \frac{16}{63}=0.254\\\\ c)\ \frac{1}{63}=0.016

C.Scoring more than 75 touchdowns is unusual.